Design a socket and spigot type cotter joint to withstand a reciprocating load of 50 kN.
Material: MS; Allowable tensile stress = 80 N/mm2
; Allowable compressive stress = 100
N/mm2
;Allowable shear stress = 50 N/mm2
.
1
Expert's answer
2020-10-08T15:11:10-0400

Load=50 kN, shear stress= "\tau" =50 "\frac{N}{mm^2}" , Allowable tensile stress="\sigma_t=" 80 "\frac{N}{mm^2}"

Allowable compressive load ="\sigma_c= 100 \frac{N}{mm^2}"

(i) For Diameter of rod


"P=\frac{\pi}{4}\times d^2\times \sigma_t"


"50\times1000=\frac{\pi}{4}\times d^2\times 80"

"d=28.21 mm" so we take dia as 30 mm

Failure of spigot in tension

"P=\frac{\pi}{4}\times d^2\times \sigma_t- dt \sigma_t, t= d/4"

"50000=\frac{\pi}{4}\times d^2\times 80- d^2\times 20"

d=34.179 mm so we take dia as 35 mm

(ii) Failure of spigot end in shear

"P=2\times a\times d\times \tau"

"50000=2\times a\times 35\times 50"

a=14.285 mm

Failure of spigot collar in shear

"P=\pi\times d\times t_1 \times \tau"

"50000=3.142\times 35 \times t_1 \times 50"

"t_1=" 9.09 mm

(iii) Failure of socket in tension

P="\frac{\pi}{4}(d_1^2-d_2^2)-(d_1-d_2)\times t\times \sigma_t"

"50000=\frac{3.14}{4}(d_1^2-35^2)-(d_1-35)\times 10\times 80"

"50000=0.785 d_1^2-961.625-800d_1+2800"

on solving this quadratic we get

d1=39.44 so we take diameter as 40 mm